OpenStudy (anonymous):

what is the average of these numbers?: 66,68,69,70,71,72,74,76,78,82,85

OpenStudy (anonymous):

\[\frac{66+68+69+70+71+72+74+76+78+82+85}{11}=\frac{811}{11}=73.7273 \]